Calculate the derivative indicated.
d2ydx2∣∣∣x=8 where y=91x3+7x2


Answer:
1
Expert's answer
2020-11-30T19:16:27-0500

"y'=91\cdot3x^2+7\cdot2x=273x^2+14x"

"\frac{d^2y}{dx^2}=y''=273\cdot2x+14=546x+14"

"\frac{d^2y}{dx^2}|_{x=8}=546\cdot8+14=4382"


Answer: 4382.